OpenStudy (anonymous):

is (6,-6) a solution of 2y-4x=-2

OpenStudy (saifoo.khan):

PLug those!

OpenStudy (gw2011):

2y-4x=-2 Substitute the values into this equation to get: 2(-6)-4(6)=-12-24=-36 Therefore, 2y-4x does not equal -2 given those values